1. In your work for a national health organization, you are asked to monitor the amount of sodium in a
certain brand of cereal. You find that a random sample of 82 cereal servings has a mean sodium content
of 232 milligrams with a standard deviation of 10 milligrams. At α = 0.01 , can you conclude that the
mean sodium content per serving of cereal is more than 230 milligrams?
(a) (DESIGN) State your Hypothesis
Ho: μ ≤ 230
Ha: μ > 230
I need help figuring how to find the P value
Answer #1

Answer

given that

test statistic t =

degree of freedom = n-1 = 82 - 1= 81

using t distribution, check df value in the top row and t value in the column, then selecting the intersecting cell, we get

p value = 0.0369

p value is greater than 0.01 significance level, so we failed to reject the null hypothesis